# Looking into the Jet Cone of the Neutrino-Associated Very High Energy Blazar PKS 1424+240

### v1

@@ 4 Emission of high energy photons and neutrinos: relativistic beaming
We note that the explanation of the VHE gamma-ray flux within a leptohadronic model of M. Cerruti et al. (2017) requires the Doppler factor of $\sim 30$, consistent with our observations. The expected associated neutrino flux is of order of that inferred from IceCube observations. The neutrino production in the blazar radio core may proceed through the $p\gamma$ mechanism, adopted by A. V. Plavin et al. (2021); O. Kalashev et al. (2023) to explain TeV to PeV neutrino production. Energetic protons, required for the neutrino and, in the hadronic scenarios, gamma-ray production, may be accelerated either in the vicinity of the central black hole (K. Ptitsyna & A. Neronov, 2016) or at the border surface between the fast spine and the slower sheath in the jet (e.g., F. Tavecchio & G. Ghisellini, 2015). In any case, the specific geometry of the jet in PKS 1424+240, which is closely aligned to the line-of-sight, provides for an additional Doppler enhancement of its multi-wavelength and multi-messenger fluxes. This boosting makes the object persistently bright, maintaining a high average flux — placing it among the top 1 % of gamma-ray sources (J. Ballet et al., 2023), and the brightest blazar in terms of high-energy neutrino emission (R. Abbasi et al., 2022). Future observations of this and other VHE emitting blazars are needed to pave the way to a quantitative model of neutrino production in jets, and to better understand the role of beaming in the gamma-ray emission.
@@ 5 Summary
Supermassive black holes at the heart of distant active galaxies are believed to be among the most powerful particle accelerators in the Universe. They produce VHE photons and serve as possible sources of cosmic neutrinos. The mechanism responsible for accelerating massive particles to relativistic energies remains an open question in modern astrophysics, with powerful extragalactic jets being a key candidate. A longstanding discrepancy—known as the Doppler factor crisis—arises from the mismatch between low apparent plasma speeds observed with VLBI and the highly relativistic flows required to explain the emission. We use 16 years of 2 cm VLBI observations to uncover properties of the parsec-scale jet of the blazar PKS 1424+240, which is exceptional in its VHE gamma-ray and probable neutrino emissions. A significant toroidal magnetic-field is detected by a clear signature in linear polarization, revealing a current-carrying jet flowing almost directly toward us. We determine that the jet is observed within its cone, at a viewing angle below half a degree. This geometry maximizes relativistic boosting while minimizing apparent speed, thereby resolving the Doppler factor crisis and suggesting extreme relativistic beaming for this VHE photon and neutrino emitting blazar. Our findings for PKS 1424+240 pave the way towards better understanding of high-energy photon and neutrino emission from blazar jets.
@@
Observing jets within their plasma cone provides a unique opportunity to determine their intrinsic parameters and reconstruct the projected magnetic field structure. Moreover, such jets may represent a distinct class of AGN with both enhanced VHE gamma-ray and neutrino emission. From Monte Carlo simulations of the complete, VLBI-selected MOJAVE 1.5JyQC sample, we estimate that only a few percent of the jets are viewed within a degree to our line of sight (M. L. Lister et al., 2019). The results of a full MOJAVE sample analysis will be presented separately by Pushkarev et al. (in prep.).
@@
We thank Andrei Lobanov, Yannis Liodakis, and Tigran Arshakian for productive discussions and comments on the manuscript. This work is part of the MuSES project which has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No 101142396). The rotation measure analysis by JDL is supported by the M2FINDERS project, which has received funding from the ERC under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101018682). AVP is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273). The work of SVT and ABP is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research made use of the data from the MOJAVE database maintained by the MOJAVE team (M. L. Lister et al., 2018).
Appendix A Supplementary Figures
@@
Figure 4: 1.6 GHz VLBA image of PKS 1424+240, epoch 9 May 2017. A circular restoring beam is shown in the left bottom corner at the FWHM level of 8.17 mas. It supports the viewing angle being less than a half-opening angle even at large (kiloparsec, deprojected) spatial scales probed by the low-frequency observations. Stokes I image FITS file is published electronically as a paper supplementary material.

### v2

@@ 4 Emission of high energy photons and neutrinos: relativistic beaming
We note that the explanation of the VHE gamma-ray flux within a leptohadronic model of Cerruti et al. (2017) requires the Doppler factor of $\sim 30$, consistent with our observations. The expected associated neutrino flux is of order of that inferred from IceCube observations. The neutrino production in the blazar radio core may proceed through the $p\gamma$ mechanism, adopted by Plavin et al. (2021) and Kalashev et al. (2023) to explain TeV to PeV neutrino production. Energetic protons, required for the neutrino and, in the hadronic scenarios, gamma-ray production, may be accelerated either in the vicinity of the central black hole (Ptitsyna & Neronov, 2016) or at the border surface between the fast spine and the slower sheath in the jet (e.g., Tavecchio & Ghisellini, 2015). In any case, the specific geometry of the jet in PKS 1424+240, which is closely aligned to the line of sight, provides an additional Doppler enhancement of its multi-wavelength and multi-messenger fluxes. This boosting makes the object persistently bright, maintaining a high average flux — placing it among the top 1% of gamma-ray sources (Ballet et al., 2023), and the brightest blazar in terms of high-energy neutrino emission (Abbasi et al., 2022). Future observations of this and other VHE emitting blazars are needed to pave the way to a quantitative model of neutrino production in jets, and to better understand the role of beaming in the gamma-ray emission.
@@ 5 Summary
Supermassive black holes at the heart of distant active galaxies are believed to be among the most powerful particle accelerators in the Universe. They produce VHE photons and serve as possible sources of cosmic neutrinos. The mechanism responsible for accelerating massive particles to relativistic energies remains an open question in modern astrophysics, with powerful extragalactic jets being a key candidate. A longstanding discrepancy — known as the Doppler factor crisis — arises from the mismatch between low apparent plasma speeds observed with VLBI and the highly relativistic jet required to explain the emission. We use 16 years of 2 cm VLBI observations to uncover properties of the parsec-scale jet of the blazar PKS 1424+240, which is exceptional in its VHE gamma-ray and probable neutrino emissions. A significant toroidal magnetic field is detected by a clear signature in linear polarization, revealing a current-carrying jet flowing almost directly toward us. We determine that the jet is observed within its cone, at a viewing angle below half a degree. This geometry maximizes relativistic boosting while minimizing apparent speed, thereby resolving the Doppler factor crisis and suggesting extreme relativistic beaming for this VHE photon and neutrino-emitting blazar. Our findings for PKS 1424+240 pave the way towards better understanding of high-energy photon and neutrino emission from blazar jets.
@@
Observing jets within their plasma cone provides a unique opportunity to determine their intrinsic parameters and reconstruct the projected magnetic field structure. Moreover, such jets may represent a distinct class of AGN with both enhanced VHE gamma-ray and neutrino emission. From Monte Carlo simulations of the complete, VLBI-selected MOJAVE 1.5JyQC sample, we estimate that only a few percent of the jets are viewed within a degree to our line of sight (Lister et al., 2019). The results of a full MOJAVE sample analysis will be presented separately by Pushkarev et al. (in prep.).
6 Data availability
Stacked Stokes I, Q, U 15 GHz FITS images (Figure 2) as well as Stokes I 1.6 GHz FITS image (Figure 4) are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
Acknowledgements.
@@
We thank Andrei Lobanov, Yannis Liodakis, and Tigran Arshakian for productive discussions and comments on the manuscript. This work is part of the MuSES project which has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No 101142396). The rotation measure analysis by JDL is supported by the M2FINDERS project, which has received funding from the ERC under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101018682). AVP is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273). The work of SVT is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research made use of the data from the MOJAVE database maintained by the MOJAVE team (Lister et al., 2018).
Appendix A Supporting observational material
A.1 VLBA data at 1.6 GHz
We have processed 1.6 GHz VLBA observations of PKS 1424+240, epoch 9 May 2017, experiment code BS257A. At ten times lower frequency, the VLBA is substantially more sensitive to distant optically thin jet emission and probes angular scales ten times larger than our MOJAVE stacked image. With a total on-source time of about 3 hr, these single epoch low-frequency observations reach a high dynamic range of about 7,000:1 and confirm the emission around the VLBI core, including the North-West direction (Figure 4). We note that our stacked image at 24 GHz also detects this emission (Figure 5).
A.2 Rotation Measure at 15, 24, and 43 GHz
As RM data is not available for all 42 epochs used in Figure 2 and individual epochs do not have RM data for the full jet region, we cannot apply a precise correction for the Faraday effect on EVPA. The median magnitude of RM for PKS 1424+240 corresponds to a EVPA rotation of $\sim 6^{\circ}$, which is comparable to our EVPA errors from calibration and imaging.
The low magnitude and variability of RM especially close to the core of PKS 1424+240 are indicative of either a weak magnetic field or low thermal electron density along the line of sight. We may be seeing the ‘sweeping away’ of the magneto-ionic medium, this would result in a reduction in the magnitude of RM. This effect would be more prominent within the center of the jet, while the edges would be less affected, which is supported by the magnitude and variability of RM being typical of other sources for the jet region of PKS 1424+240.
@@ A.2 Rotation Measure at 15, 24, and 43 GHz
Figure 4: 1.6 GHz VLBA image of PKS 1424+240, epoch 9 May 2017. A circular restoring beam is shown in the left bottom corner at the FWHM level of 8.17 mas. The image supports the viewing angle being less than a half-opening angle even at large (kiloparsec, deprojected) spatial scales probed by the low-frequency observations. Stokes I image FITS file is is available at the CDS.